### v2

# When Do Credal Sets Stabilize? Fixed-Point Theorems for Credal Set Updates

Abstract.
Many machine learning algorithms rely on iterative updates of uncertainty representations, ranging from variational inference and expectation-maximization, to reinforcement learning, continual learning, and multi-agent learning. In the presence of imprecision and ambiguity, credal sets—closed, convex sets of probability distributions—have emerged as a popular framework for representing imprecise probabilistic beliefs. Under such imprecision, many learning problems in imprecise probabilistic machine learning (IPML) may be viewed as processes involving successive applications of update rules on credal sets. This naturally raises the question of whether this iterative process converges to stable fixed points—or, more generally, under what conditions on the updating mechanism such fixed points exist, and whether they can be attained. We provide the first analysis of this problem, and illustrate our findings using Credal Bayesian Deep Learning as a concrete example. Our work demonstrates that incorporating imprecision into the learning process not only enriches the representation of uncertainty, but also reveals structural conditions under which stability emerges, thereby offering new insights into the dynamics of iterative learning under imprecision.
Key words and phrases:
Fixed Point Theory; Credal Sets; Probabilistic Machine Learning
2020 Mathematics Subject Classification:
Primary: 54H25; Secondary: 68T05; 68T37
1. Introduction
In learning problems, a fixed point refers to a solution that remains unchanged when a specific transformation or algorithm is applied repeatedly. A fixed point’s existence often acts as a guarantee of stability for an iterative process, meaning the system’s dynamics will eventually converge to a stable solution (Banach,, 1922; Ortega and Rheinboldt,, 2000). Fixed point theorems have been used in proving that a learning or optimization problem has a solution—e.g. an equilibrium point, stable model, or policy (Banach,, 1922; Kakutani,, 1941)—showing that an iterative learning rule—e.g. policy update, message passing, and Expectation Maximization (EM) step—converges to a stable point (Dempster et al.,, 1977; Bertsekas and Tsitsiklis,, 1996), and characterizing Nash equilibria or best-response dynamics of agents’ strategies in multi-agent learning (Nash,, 1950; Fudenberg and Tirole,, 1991). Recently, fixed point analysis has become a valuable tool for understanding the dynamics that arise from the interaction between human and AI systems—such as phenomena like model collapse (Shumailov et al.,, 2024) and performativity (Perdomo et al.,, 2020; Hardt and Mendler-Dünner,, 2025) that emerge from these complex interactions—as well as reciprocal learning where multiple agents, models, or systems adapt to each other’s behavior in a mutually dependent way (Rodemann et al.,, 2024; Rodemann and Bailie,, 2025).
Classical fixed point analysis, though widely applied, often falls short of capturing the intrinsic ambiguity in learning problems. In the domain of continual learning, where models are trained iteratively on a stream of data, the learning dynamics frequently exhibit instability (Kirkpatrick et al.,, 2017; Parisi et al.,, 2019). Empirical observations reveal that the training process often fails to converge to a single, stable solution. This lack of convergence manifests as significant fluctuations in the model’s parameters, which can substantially degrade performance on previously learned tasks. This phenomenon, widely known as catastrophic forgetting, represents a key challenge in developing robust continual learning systems. Furthermore, a fixed point for iterative update rules may not be attainable when the underlying data sources contain inherent conflicts (Arrow,, 1950; Walley,, 1991; Genest and Zidek,, 1986). These conflicts, which can arise from disparate sources like heterogeneous datasets or conflicting individual opinions, introduce a dynamic instability that prevents convergence to a stable solution.
Inspired by these challenges, this paper develops fixed point theorems for update rules on credal sets—closed, convex sets of probability distributions (Levi,, 1980; Walley,, 2000)—which remain comparatively understudied relative to fixed point analysis on the space of single probability distributions. Update rules on credal sets play a central role in imprecise probabilistic machine learning (IPML), a growing field that leverages imprecise probability (IP) theory to enhance robustness, trustworthiness, and safety in machine learning. Classical examples include the generalized Bayes rule (Walley,, 1991), Dempster’s rule (Dempster,, 1967; Shafer,, 1976), and the geometric rule (Suppes and Zanotti,, 1977), all of which operate directly on credal sets. More recently, modern approaches such as Credal Bayesian Deep Learning (CBDL) (Caprio et al., 2024a, ) and Imprecise Bayesian Continual Learning (IBCL) (Lu et al.,, 2025) have applied these rules in the context of Bayesian deep learning (BDL) and continual learning (CL). Their historical significance, combined with their increasing adoption in contemporary machine learning, underscores the importance of establishing fundamental conditions under which the stable solutions exist.
Our contributions. We establish structural conditions under which update rules for credal sets attain stability. Our main contributions are as follows:
  • Our first result, Theorem 1, gives the minimal conditions under which an updating technique $f$ for credal sets admits a fixed point. We also show that the conditions are truly minimal, and study when the Credal Bayesian Deep Learning (CBDL) paradigm (Caprio et al., 2024a, ) satisfies them.
  • Next, we inspect when the fixed point found in Theorem 1 is unique, and when the sequence of successive $f$-updating of a credal set converges to such a unique fixed point. Continuing our running example, we provide the conditions under which CBDL exhibits a unique fixed point, and convergence to it.
  • Finally, we provide an inner and outer approximation for a sequence of credal sets updated using different updating functions. This version of the squeeze theorem has promising implications for IPML-backed continual and active learning.
To demonstrate our findings empirically, we conduct a simple synthetic experiment on finitely generated credal sets, that is, credal sets having finitely many extreme elements, viz. elements that cannot be written as a convex combination of one another.
Given the page limitations, we provide the necessary mathematical background—namely the definitions of compact metric space, weak topology (in this paper, we call weak-convergence the convergence in expectation of all bounded continuous functions), dominating measure, and likelihood kernel, as well as the statements of Kakutani’s, Edelstein’s, and Boyd-Wong’s fixed point theorems—in Appendix A, and the proofs of our results in Appendix B.
2. Preliminary and Related Work
2.1. Fixed Point Theorems
Fixed point theory (Granas et al.,, 2003) studies the existence, uniqueness, and qualitative properties of solutions to equations of the form $f(x)=x$, for a given mapping $f$. Its applications are widespread across disciplines. In economics, fixed point results underpin the existence of equilibria in game theory and general equilibrium models (Scarf,, 1983). In physics, they are used to analyze the long-term behavior of dynamical systems and phase transitions (Dancer and Hess,, 1991). In computer science and logic, fixed points provide the foundation for induction principles and recursive definitions, ensuring that self-referential constructions are mathematically well grounded (Abel,, 2012).
In probability theory and statistics, fixed point theorems are crucial for establishing the existence of invariant probability measures. Examples include stationary distributions of Markov chains and equilibrium laws in mean-field models. Here, the space of probability measures provides a natural setting in which classical results such as the Banach’s (Banach,, 1922), Schauder’s (Schauder,, 1930), and Kakutani’s (Kakutani,, 1941) fixed point theorems can be applied to prove both existence and uniqueness of such invariant distributions.
However, when viewed from a probabilistic perspective, these classical results are typically formulated for mappings acting on precise probability measures. Their applicability becomes less clear when considering mappings on the space of credal sets (Levi,, 1980)—that is, nonempty, weak-closed, convex subsets of probability measures. Such sets extend beyond the finite dimensional setting of probability vectors and introduce additional structural complexity. Investigating how fixed point results can be generalized to this richer setting is the focus of Section 3. Pursuing this direction promises to provide a rigorous theoretical foundation for imprecise probabilistic machine learning, introduced in the following subsection, where credal sets play a central role.
It is worth noting that the field of imprecise stochastic processes—especially the work on the convergence of imprecise Markov chains (de Cooman et al.,, 2008, 2009; Škulj and Hable,, 2009)—addresses questions that are similar in spirit to those studied in this paper, although typically restricted to (at most) countable state spaces. In particular, their results rely on the Perron-Frobenius theorem, which we conjecture could also be derived using our fixed point methods.
2.2. Imprecise Probabilistic Machine Learning
Imprecise Probabilistic Machine Learning (IPML) is an emerging field that seeks to integrate the rigorous theory of imprecise probabilities (Walley,, 1991) into probabilistic machine learning. Its ultimate goal is to be precise about imprecision, thereby enabling more robust and trustworthy models and inferences. Among the various representations of imprecision, the credal set is arguably the most central object of study. In recent years, the incorporation of credal sets into machine learning has led to significant developments across a wide range of areas, including classification (Caprio et al., 2024b, ), domain generalization (Singh et al.,, 2024), hypothesis testing (Chau et al., 2025b, ; Jürgens et al.,, 2025), scoring rules (Fröhlich and Williamson,, 2024; Singh et al.,, 2025), conformal prediction (Caprio et al.,, 2025; Caprio,, 2025), computer vision (Cuzzolin and Frezza,, 1999; Giunchiglia et al.,, 2023), probabilistic programming (Liell-Cock and Staton,, 2025), explainability (Chau et al.,, 2023; Utkin et al.,, 2025; Mohammadi et al.,, 2026), neural networks (Caprio et al., 2024a, ; Wang et al.,, 2024, 2025), learning theory (Caprio et al., 2024c, ), causal inference (Cozman,, 2000; Zaffalon et al.,, 2023), active and continual learning (Dutta et al.,, 2024; Lu et al.,, 2025), and probability metrics (Chau et al., 2025a, ), among others.
However, to the best of our knowledge, there has been no prior work studying fixed point theorems for mappings acting on credal sets. Filling this gap is the main focus of our contribution.
Figure 1. An illustration of Theorem 1 with sequences of credal sets defined in a 2d probability simplex. As we apply our Hausdorff continuous update rule $f$ to a credal set repeatedly, it eventually arrives at a fixed “point”.
3. Main Results
Let $\mathcal{X}$ be a measurable space of interest, and call $\Delta_{\mathcal{X}}$ the space of countably additive probability measures on $\mathcal{X}$. While usually in Imprecise Probability theory scholars work with finitely additive probabilities, in this section we focus on the countably additive case to avoid complications. The finitely additive case is studied in Appendix D. Let $\mathscr{C}\subset 2^{\Delta_{\mathcal{X}}}$ be the space of nonempty weak-closed and convex subsets of $\Delta_{\mathcal{X}}$. Let $\mathcal{P}$ be a generic credal set on $\mathcal{X}$ (Levi,, 1980), so that $\mathcal{P}\in\mathscr{C}$.
3.1. Existence of a Fixed Point
Suppose now that we wish to model an updating mechanism of the agent’s beliefs, that are encapsulated in $\mathcal{P}$. We show that if such an updating can be expressed as a sufficiently well-behaved function, then our updating method enjoys desirable fixed point properties.
Recall that a function $f:\mathscr{C}\rightarrow\mathscr{C}$ is Hausdorff continuous if, for a net $(\mathcal{P}_{\gamma})$ in $\mathscr{C}$, $\mathcal{P}_{\gamma}\xrightarrow[]{d_{H}}\mathcal{P}$ implies $f(\mathcal{P}_{\gamma})\xrightarrow[]{d_{H}}f(\mathcal{P})$, where $d_{H}$ denotes the Hausdorff metric. The latter is defined as $d_{H}(\mathcal{P},\mathcal{Q})\coloneqq\max\{\sup_{P\in\mathcal{P}}\rho(P,\mathcal{Q}),\sup_{Q\in\mathcal{Q}}\rho(\mathcal{P},Q)\}$, where $\rho(P,\mathcal{Q})\coloneqq\inf_{Q\in\mathcal{Q}}\rho(P,Q)$ and similarly for $\rho(\mathcal{P},Q)$, and $\rho$ is any metric that metrizes the weak topology, like the Prokhorov metric (Billingsley,, 2013, Theorem 6.8).
Theorem 1 (Fixed Point Theorem for Credal Updating).
Let $\mathcal{X}=(\mathcal{X},d)$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is continuous in the Hausdorff topology, then the set of fixed points of $f$ is compact and nonempty.
Theorem 1, proved in Appendix B.1, tells us that, if our space $\mathcal{X}$ of interest has a sufficiently well-behaved topological structure, then a continuous updating function is guaranteed to have (at least) a fixed point. Theorem 1 bears similarities to a version of Kakutani’s Fixed Point Theorem, presented in Appendix A.5.
Notice that the assumption in Theorem 1 that $\mathcal{X}$ is compact is not overly strong: For example, the space of colored images $0,1]^{w\times h\times 3}$ (where $w$ denotes the images’ width, $h$ their height, and $3$ stands for the RGB option) is compact. In practice, if $\mathcal{X}$ is not compact (in the topology compatible with the metric $d$ that we choose), we can always either change the metric (and hence also its compatible topology), or consider the compactification of $\mathcal{X}$. [^1] In that case, the weak topology on $\Delta_{\mathcal{X}}$ is understood with respect to the new topology on $\mathcal{X}$, and may differ from the original one. The real heavy lifting is done by the updating function $f$ being continuous (and assumed to output credal sets). This key property cannot be arbitrarily modified, as shown by the following example.
Example 1 (On the Importance of a Continuous Updating Rule).
Let $\mathcal{X}=\{0,1\}$, so $\Delta_{\mathcal{X}}\cong[0,1]$ via $P_{\lambda}(\{1\})=\lambda$, $P_{\lambda}(\{0\})=1-\lambda$. Let $\mathscr{C}$ be the family of all nonempty closed convex subsets of $\Delta_{\mathcal{X}}$, i.e., all closed intervals $\lambda_{1},\lambda_{2}]\subseteq[0,1]$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$. Fix $\delta\in(0,\tfrac{1}{2})$ and define $f:\mathscr{C}\to\mathscr{C}$ by
$f([\lambda_{1},\lambda_{2}])\coloneqq\begin{cases}[\lambda_{1}+\delta,\lambda_{2}+\delta],&\text{if }\lambda_{2}<1-\delta,\\[4.0pt] \big([\lambda_{1}-\delta,\lambda_{2}-\delta]\cap[0,1]\big),&\text{if }\lambda_{2}\geq 1-\delta.\end{cases}$
Then,
  • Well-definedness. If $\lambda_{2}<1-\delta$, then $\lambda_{2}+\delta<1$ and $\lambda_{1}+\delta\leq\lambda_{2}+\delta$, so the first case stays in $[0,1]$. If $\lambda_{2}\geq 1-\delta$, subtracting $\delta$ may push the left endpoint below 0, hence the intersection with $[0,1]$ yields a (nonempty) closed interval; Thus $f(\mathscr{C})\subseteq\mathscr{C}$.
  • Discontinuity (Hausdorff). Fix any $a\in[0,1-\delta]$ and let $I_{n}=[a,1-\delta-1/n]\to I=[a,1-\delta]$ in $d_{H}$. Then $f(I_{n})=[a+\delta,1-1/n]\to[a+\delta,1]$, while $f(I)=[\max\{0,a-\delta\},1-2\delta]$. Hence $d_{H}\big(f(I_{n}),f(I)\big)\geq\delta$, so $f$ is not Hausdorff-continuous at $\lambda_{2}=1-\delta$.
  • No fixed points. If $\lambda_{2}<1-\delta$, then $f$ shifts the interval right by $\delta$, so equality would force $\delta=0$, a contradiction. If $\lambda_{2}\geq 1-\delta$, then $f([\lambda_{1},\lambda_{2}])=[\max\{0,\lambda_{1}-\delta\},\lambda_{2}-\delta]$, so equality would require $\lambda_{2}=\lambda_{2}-\delta$, again a contradiction. Thus, the set $\mathrm{Fix}(f)$ of fixed points of $f$ is empty.
This shows that without Hausdorff continuity, even on a compact credal hyperspace, fixed points for the update rule need not exist.
In applications, one rarely verifies Hausdorff continuity of $f:\mathscr{C}\to\mathscr{C}$ from first principles. Instead, one typically exploits structural properties of the credal sets and of the updating rule, or resorts to numerical diagnostics. We briefly outline three practically relevant situations, followed by an example based on Credal Bayesian Deep Learning, an IPML updating technique based on a generalization of Bayes’ updating.
(1) Finite-dimensional/polyhedral representations. Let $\Delta_{\mathcal{X}}$ be a finite-dimensional simplex, and let credal sets $\mathcal{P}\in\mathscr{C}$ be represented as polytopes, for instance $\mathcal{P}(\eta)=\{p\in\Delta_{\mathcal{X}}:A(\eta)p\leq b(\eta)\}$ or $\mathcal{P}(\eta)=\operatorname{CH}\{p_{1}(\eta),\dots,p_{m}(\eta)\}$, where $\text{CH}(\cdot)$ denotes the convex hull operator, for some parameter $\eta$ encoding vertices or constraints. If the updater is built from operations such as pointwise application of a continuous map $T:\Delta_{\mathcal{X}}\to\Delta_{\mathcal{X}}$, followed by convex hull, $f(\mathcal{P})=\operatorname{CH}\{T(p):p\in\mathcal{P}\}$, or intersection with families of half-spaces whose coefficients depend continuously on $\eta$, then the induced map $\mathcal{P}(\eta)\mapsto f(\mathcal{P}(\eta))$ is continuous with respect to the Hausdorff metric $d_{H}$. Intuitively, small perturbations of the vertices or constraints of $\mathcal{P}$ lead to small perturbations of the vertices or constraints of $f(\mathcal{P})$, and hence to small changes in $f(\mathcal{P})$ in the Hausdorff topology.
(2) Optimization-based updating rules. A second possible pattern is that the updated credal set is defined as the solution set of an optimization problem,
$f(\mathcal{P})=\operatorname*{argmin}_{Q\in\mathcal{A}(\mathcal{P})}\mathfrak{L}(Q),$
where $\mathfrak{L}$ is a loss or divergence functional and $\mathcal{A}(\mathcal{P})$ is a feasible region derived from the initial credal set $\mathcal{P}$ and the observed data. Under standard stability conditions from parametric optimization, for example (i) $\mathfrak{L}$ is continuous (or even strictly convex) in $Q$; (ii) the feasible set $\mathcal{A}(\mathcal{P})$ is nonempty, compact, and depends continuously on $\mathcal{P}$ (e.g., via constraints whose coefficients vary continuously with parameters encoding $\mathcal{P}$); (iii) minimizers are unique or vary continuously as these parameters change, the solution mapping $\mathcal{P}\mapsto f(\mathcal{P})$ is continuous in the Hausdorff metric. Thus, optimization-based updating techniques can inherit Hausdorff continuity from the continuity of their objective and constraints.
(3) Heuristic and numerical diagnostics. Even when an analytic verification is cumbersome, practitioners can still assess the behavior of $f$ heuristically. A simple strategy is:
  1. (1)
    Parametrize credal sets $\mathcal{P}_{\eta}\in\mathscr{C}$ by a finite-dimensional parameter $\eta$ (e.g. bounds, vertices, or hyperparameters) in a normed parameter space.
  2. (2)
    For small perturbations $\kappa\eta$, compute $f(\mathcal{P}_{\eta})$ and $f(\mathcal{P}_{\eta+\kappa\eta})$.
  3. (3)
    Approximate the Hausdorff distance $d_{H}\bigl(f(\mathcal{P}_{\eta}),f(\mathcal{P}_{\eta+\kappa\eta})\bigr)$ by sampling a finite set of distributions from each credal set and computing pairwise distances.
If $d_{H}(f(\mathcal{P}_{\eta}),f(\mathcal{P}_{\eta+\kappa\eta}))$ remains small and scales with $\|\kappa\eta\|$ across a variety of directions in parameter space (and, if relevant, across small perturbations of the data), this provides empirical evidence that the updater behaves in a Hausdorff-continuous way. Conversely, large, irregular jumps in these approximate distances typically signal violations of continuity (for instance caused by hard thresholds, selection of a single “best” model, or conditioning on near-zero likelihood events).
These observations indicate that, in many imprecise probabilistic ML settings of interest, Hausdorff continuity of $f$ is either guaranteed by the way the updater is constructed (polyhedral or optimization-based rules) or can at least be inspected numerically. Let us now turn our attention to Credal Bayesian Deep Learning.
3.1.1. Credal Bayesian Deep Learning (CBDL)
In Caprio et al., 2024a , the authors propose Credal Bayesian Deep Learning (CBDL), a novel framework for Bayesian deep learning. At training time, the practitioner specifies prior and likelihood Finitely Generated Credal Sets (FGCSs): The prior FGCS on the network parameters and the likelihood FGCS to possibly encode different network architectures. Then, the posterior credal set is derived by computing the Bayesian updating of all possible pairs of priors and likelihoods that are extrema of their respective credal sets. This can be formalized as follows.
Let $\mathcal{P}\in\mathscr{C}\subset 2^{\Delta_{\Theta}}$ be our prior FGCS, where $\Theta$ is our parameter space; Denote the collection of its extreme elements as $\text{ex}\mathcal{P}$, and its members by $P^{\text{ex}}$. Let $\mathcal{L}\subset 2^{\Delta_{\mathcal{Y}}}$ be our likelihood FGCS (where $\mathcal{Y}$ is the measurable sample space), and let $\text{ex}\mathcal{L}=\{L_{1},\ldots,L_{K}\}$, $K\in\mathbb{N}$.
Let $\mu$ be a $\sigma$-finite dominating measure (see Appendix A.3) for $\text{ex}\mathcal{P}$ on $\Theta$. For $P^{\mathrm{ex}}$ with density $p^{\mathrm{ex}}$ with respect to $\mu$ and likelihood kernel $\ell_{k}(E\mid\theta)$ (see Appendix A.4), set $P^{\mathrm{ex}}(A)=\int_{A}p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)$, $L_{k}(E)=\int_{\Theta}\ell_{k}(E\mid\theta)p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)$, and define the prior-conditional evidence
$L_{k}(E\mid A)=\begin{cases}\displaystyle\frac{\int_{A}\ell_{k}(E\mid\theta)p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)}{P^{\mathrm{ex}}(A)},&\text{if }P^{\mathrm{ex}}(A)>0,\\[8.0pt] 0,&\text{if }P^{\mathrm{ex}}(A)=0.\end{cases}$
Recall that a correspondence $\varphi:\Delta_{\Theta}\rightrightarrows\Delta_{\Theta}$ assigns to some $P\in\Delta_{\Theta}$ a subset $\varphi(P)\subset\Delta_{\Theta}$, and that $\varphi(A)\coloneqq\cup_{P\in A}\varphi(P)$, for all $A\subset\Delta_{\Theta}$ (Aliprantis and Border,, 2006, Section 17.1). The updating procedure prescribed by CBDL for $\mathcal{P}$, can be written as a function $f:\mathscr{C}\rightarrow\mathscr{C}$ such that
(1) $\mathcal{P}\mapsto f(\mathcal{P})\coloneqq\text{CH}\left(\varphi(\text{ex}\mathcal{P})\right),$
where $\varphi:\Delta_{\Theta}\rightrightarrows\Delta_{\Theta}$ is a correspondence $\text{ex}\mathcal{P}\ni P^{\text{ex}}\mapsto\varphi(P^{\text{ex}})\coloneqq\{g_{1}(P^{\text{ex}}),\ldots,g_{K}(P^{\text{ex}})\},$ where $K\in\mathbb{N}$, and for $k\in\{1,\dots,K\}$,
(2) $g_{k}(P^{\text{ex}})(A)\coloneqq\frac{P^{\text{ex}}(A)L_{k}(E\mid A)}{L_{k}(E)}$
for measurable sets $A\subset\Theta$, and measurable $E\subset\mathcal{Y}$ such that $L_{k}(E)\neq 0$. Here, of course, the function $g_{k}$ computes the Bayesian updating of $P^{\text{ex}}$ via likelihood $L_{k}\in\text{ex}\mathcal{L}$. The following is proven in Appendix B.2.
Proposition 2.
Fix the measurable evidence set $E\subset\mathcal{Y}$. If the following three hold,
  1. (I)
    The parameter space $\Theta$ is compact metric,
  2. (II)
    For all $P^{\text{ex}}\in\text{ex}\mathcal{P}$, $p^{\text{ex}}\in\mathscr{L}^{1}(\Theta,\mu)$, that is, $\int_{\Theta}|p^{\text{ex}}(\theta)|\mu(\text{d}\theta)<\infty.$ Futhermore, for all $L_{k}\in\text{ex}\mathcal{L}$, $\ell_{k}(E\mid\cdot)$ is continuous,
  3. (III)
    For all $L_{k}\in\text{ex}\mathcal{L}$, $\ell_{k}(E\mid\theta)>0$, for all $\theta\in\Theta$,
then the function $f$ is Hausdorff-continuous.
We can then apply Theorem 1 to show that the CBDL updating technique for a prior FGCS $\mathcal{P}$ defined in Caprio et al., 2024a (, Algorithm 1) has (at least) a fixed point. A similar result holds also for imprecise Bayesian continual learning (Lu et al.,, 2025), a continual learning (CL) method in which successive updates of credal sets are used to address CL under specific trade-offs.
3.2. Uniqueness of a Fixed Point
The next corollary shows that if $f$ satisfies a strong form of non-expansiveness—namely it is strictly distance-decreasing in the Hausdorff metric—then its fixed point is unique. It is proved in Appendix B.3
Corollary 2.1 (Unique Fixed Point Theorem for Credal Updating).
Let $\mathcal{X}$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ satisfies
$d_{H}(f(\mathcal{P}),f(\mathcal{Q}))<d_{H}(\mathcal{P},\mathcal{Q}),\quad\forall\mathcal{P},\mathcal{Q}\in\mathscr{C}\text{, }\mathcal{P}\neq\mathcal{Q},$
then $f$ has a unique fixed point.
3.3. Convergence of the Orbit to the Unique Fixed Point
Although interesting, Corollary 2.1 is not enough to ensure that the sequence $(\mathcal{P}_{n})$ of successive updating of an initial credal set $\mathcal{P}_{0}\in\mathscr{C}$, where $\mathcal{P}_{n+1}=f(\mathcal{P}_{n})$, for all $n\in\mathbb{Z}_{+}$—called the orbit of $\mathcal{P}_{0}$ under ff—converges in the Hausdorff metric $d_{H}$ to a unique limiting credal set $\mathcal{P}_{\star}$. For this to hold, we need to require a somewhat stronger condition, and assume that $\mathcal{P}_{n}\neq\emptyset$, for all $n\in\mathbb{Z}_{+}$. We next present a generalization to credal sets of Boyd-Wong’s fixed point theorem, whose classical version is provided in Appendix A.5. There, we also provide a remark justifying our choice to generalize Boyd–Wong to study the convergence of the orbit $(\mathcal{P}_{n})$: This approach allows us to state the result in maximal generality.
Theorem 3 (Convergence of Orbit $(\mathcal{P}_{n})$).
Let $\mathcal{X}$ be a Polish space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is such that
$d_{H}(f(\mathcal{P}),f(\mathcal{Q}))\leq\psi(d_{H}(\mathcal{P},\mathcal{Q})),\quad\forall\mathcal{P},\mathcal{Q}\in\mathscr{C}\text{, }\mathcal{P}\neq\mathcal{Q},$
where $\psi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is
  1. (i)
    upper semicontinuous from the right on $\mathbb{R}_{+}$,
  2. (ii)
    such that $\psi(t)<t$, for all $t>0$,
then $f$ has a unique fixed point $\mathcal{P}_{\star}$, and, for each $\mathcal{P}_{0}\in\mathscr{C}$, the orbit $(\mathcal{P}_{n})$ of $\mathcal{P}_{0}$ under $f$ converges to $\mathcal{P}_{\star}$. That is, $\mathcal{P}_{n}\xrightarrow[n\rightarrow\infty]{d_{H}}\mathcal{P}_{\star}$.
Notice how Theorem 3, proved in Appendix B.4, requires $f$ to be a contraction, but it allows the contraction rate to vary with the distance between credal sets. For example, for $\mathcal{P}$ and $\mathcal{Q}$ that are close, i.e. for which $d_{H}(\mathcal{P},\mathcal{Q})$ is small, $f$ may contract greatly, while for $\mathcal{P}$ and $\mathcal{Q}$ that are far, i.e. for which $d_{H}(\mathcal{P},\mathcal{Q})$ is large, then $f$ may contract only slightly—but still producing two images that are strictly closer to each other. Notice also that (a) the only information from Theorem 3 about the convergence rate of $(\mathcal{P}_{n})$ is that it is sublinear and governed by $\psi$; and (b) if we require the stronger condition of $\mathcal{X}$ being compact metric, then Theorem 3 becomes a corollary of Theorem 1.
In practice, conditions (i) and (ii) in Theorem 3 can be verified by showing that $f$ is a strict contraction in the Hausdorff metric, i.e. there exists $L<1$ such that $d_{H}(f(\mathcal{P}),f(\mathcal{Q}))\leq Ld_{H}(\mathcal{P},\mathcal{Q})$ for all $\mathcal{P},\mathcal{Q}\in\mathscr{C}$, which corresponds to the special case $\psi(t)=Lt$. The structural and heuristic checks in Section 3.1 can then be adapted to assess this contraction behavior empirically.
3.3.1. CBDL, continued
Continue from the setting of Section 3.1.1, we restrict our attention to priors in $\Delta_{\Theta}^{++}$, i.e. those with densities bounded away from 0 and $\infty$ with respect to $\mu$. To see why, notice that, for any $\theta_{1}\neq\theta_{2}$, $T_{k}(\delta_{\theta_{1}})=\delta_{\theta_{1}}$ and $T_{k}(\delta_{\theta_{2}})=\delta_{\theta_{2}}$, so no metric that separates atoms can give a global contraction constant <1<1 on $\Delta_{\Theta}$. Excluding atoms, and therefore working with $\Delta_{\Theta}^{++}$, is necessary.
Proposition 4.
Fix the measurable set $E\subset\mathcal{Y}$ representing the gathered evidence. Suppose that conditions (I) and (II) in Proposition 2 still hold, but require the following strengthening of the condition in (III) for the likelihoods $L_{k}\in\text{ex}\mathcal{L}$,
  • (III’)
    For all $L_{k}\in\text{ex}\mathcal{L}$, there exist constants $0<\alpha_{k}\leq\beta_{k}<\infty$ such that $\alpha_{k}\leq\ell_{k}(E\mid\theta)\leq\beta_{k},$ for all $\theta\in\Theta$.
Then, we can find a function $\psi$ that satisfies (i)-(ii) in Theorem 3. It is
$t\mapsto\psi(t)\coloneqq\sup_{(\mathcal{P},\mathcal{Q})}\left\{d_{H}(f(\mathcal{P}),f(\mathcal{Q})):d_{H}(\mathcal{P},\mathcal{Q})\leq t\right\}.$
Proposition 4 implies that, under conditions (I), (II), and (III’), we can apply Theorem 3 to conclude that the CBDL updating technique for a prior FGCS $\mathcal{P}$ has a unique fixed point $\mathcal{P}_{\star}$, and that the sequence $(\mathcal{P}_{n})$ of the orbits of $\mathcal{P}$ under $f$ converges to $\mathcal{P}_{\star}$. In addition, the proof of Proposition 4 (which can be found in Appendix B.5) gives us a nice intuition. If two prior FGCSs are close, their updates get even closer, capturing a kind of informational convergence under CBDL and assumptions (I), (II), (III’).
Remark 1 (Convergence Rate for CBDL).
It is interesting to find the rate of convergence for the sequence $(\mathcal{P}_{n})$ in CBDL. Let us denote by $(\mathcal{P}_{n})$ the sequence of orbits of a prior FGCS $\mathcal{P}$ under $f$. Let $d_{n}\coloneqq d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})$, where $\mathcal{P}_{\star}$ is the unique fixed point of $f$, as per Theorem 3, and $d_{H}$ denotes the Hausdorff metric.
Then by Proposition 10 in Appendix B.5, the control function $\psi$ satisfies the linear bound
$\psi(t)\leq\tau t,\qquad\tau\coloneqq\max_{1\leq k\leq K}\tanh\Big(\frac{1}{4}\log\frac{\beta_{k}}{\alpha_{k}}\Big)\in(0,1).$
Hence,
$d_{n+1}=d_{H}\big(f(\mathcal{P}_{n}),f(\mathcal{P}_{\star})\big)\leq\psi\big(d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})\big)\leq\tau d_{n},$
and therefore $d_{n}\leq\tau^{n}d_{0}$ for $n\geq 0$. In particular, the orbit converges to $\mathcal{P}_{\star}$ at a geometric rate, i.e., $d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})=\mathcal{O}(\tau^{n}).$
Assumptions used. This bound relies on the standing assumptions from Section 3.3.1: Priors in $\Delta_{\Theta}^{++}$ and likelihoods satisfying (III’).
At this point, the attentive reader may have noticed that, in Proposition 4, we fixed the evidence $E\subset\mathcal{Y}$. While the result that we have proved is correct, then, it is arguably not very interesting: It is very seldom the case that we keep observing the same evidence at every step $n\in\mathbb{Z}_{+}$. To give a more compelling result, we need to consider the following modification of CBDL, that we call Pessimistic CBDL (PCBDL). Keep everything else unchanged, but, for any measurable set $A\subset\Theta$ and any $k\in\{1,\ldots,K\}$, modify (2) to
(3) $\underline{g}_{k}(P^{\text{ex}})(A)\coloneqq\inf_{E\in\mathcal{E}}\frac{P^{\text{ex}}(A)L_{k}(E\mid A)}{L_{k}(E)},$
where $\mathcal{E}\subset\mathcal{F}_{\mathcal{Y}}\setminus\{\emptyset\}$ (recall $\mathcal{F}_{\mathcal{Y}}$ is the $\sigma$-algebra on $\mathcal{Y}$), is a collection of subsets of $\mathcal{Y}$ whose elements are measurable and such that the ratio is well defined. Note that, for a given set $A\subseteq\Theta$, the event $E\in\mathcal{E}$ at which the infimum in (3) is attained may depend on $A$. Thus, the choice of $E$ used to compute $\underline{g}_{k}(P^{\text{ex}})(A)$ can vary with $A$.
Condition (3) tells us two things. First, instead of fixing $E$, we fix a class $\mathcal{E}$ of measurable sets that are associated with the experiment at hand. That is, no matter what step $n\in\mathbb{Z}_{+}$ we are considering, we know that the evidence $E$ that we may gather belongs to $\mathcal{E}$. To borrow terminology from Walley, (1991, Section 2.1.2), we say that the elements of $\mathcal{E}$ are the pragmatic possible outcomes of the experiment. For example, if we consider a coin tossing, we may posit that $\mathcal{E}=2^{\{H,T\}}\setminus\{\emptyset\}$, where $H$ denotes heads and $T$ denotes tails, instead of considering the power set of all apparently possible outcomes (Caprio and Mukherjee,, 2023, Section 3). These include e.g. coin landing on its edge, coin breaking into pieces on landing, coin disappearing down a crack in the floor.
Second, at every step $n\in\mathbb{Z}_{+}$, the previous parameter beliefs are updated, with respect to model $k$, according to the most pessimistic evidence possible (that is, the one resulting in the smallest possible updated value that $\underline{g}_{k}(P^{\text{ex}})$ attaches to the parameter event $A$), captured by the infimum over the elements of $\mathcal{E}$. At this point, a critical reader may argue that—if we specify $\mathcal{E}$ ex ante—PCBDL can be applied without collecting evidence altogether. While this is true in principle, in practice it will hardly be the case: The data gathered at step $n=1$ will inform the elicitation of $\mathcal{E}$, and if at some step $n>1$ we realize that a richer $\mathcal{E}$ is needed, we can restart the updating process, and base it on the larger evidence class.
Now we must ask ourselves whether $\underline{g}_{k}$ defined in (3) is continuous. The answer is no in general, and if we do not require some extra structure on $\mathcal{E}$, the results derived so far do not apply to PCBDL. We also have to be extra careful to ensure that $\underline{g}_{k}(P^{\text{ex}})$ is indeed a countably additive probability measure. We need the following,
  • (IV’)
    $\mathcal{Y}$ is compact metric, $\mathcal{E}\subset\mathcal{K}(\mathcal{Y})$, the collection of nonempty compact subsets of $\mathcal{Y}$, and conditions (II) and (III’) hold for all $E\in\mathcal{E}$. In addition, the evidence $E_{\star}$ that minimizes (3) is the same for all measurable $A\subset\Theta$, that is, $E_{\star}\equiv E_{\star}(A)$, for all $A$.
Note that $E_{\star}$ does not depend on $P^{\text{ex}}$. It is also easy to see that condition (IV’), and in particular $E_{\star}\equiv E_{\star}(A)$, for all $A$, is particularly strong. Under conditions (I), (II), (III’), and (IV’), the $\underline{g}_{k}$’s are continuous and the $\underline{g}_{k}(P^{\text{ex}})$’s are countably additive probabilities, so all the results we derived previously continue to hold.
More in general, our results hold for modifications to the CBDL updating mechanism that ensure continuity of (a modification of) the $g_{k}$’s, and that (a modification of) the ${g}_{k}(P^{\text{ex}})$’s are Kolmogorovian probabilities.
3.4. Limiting Behavior Under Different Updating Functions
Section 3.3.1 shows that satisfying the assumptions of Theorem 3 can sometimes be challenging. This motivates studying the “orbit” of a credal set $\mathcal{P}_{0}$ under a sequence of functions $f_{i}$, $i\in\{1,\ldots,n\}$, as $n\to\infty$, which is the focus of this section.
Suppose that we collect evidence in the form of sets $E_{1},\ldots,E_{m}\in\mathcal{F}_{\mathcal{Y}}$, $m\in\mathbb{N}_{\geq 2}$, where $\mathcal{F}_{\mathcal{Y}}$ denotes the $\sigma$-algebra of the evidence space $\mathcal{Y}$, possibly different from $\mathcal{X}$. Call $\breve{E}\coloneqq\cup_{i=1}^{m}E_{i}$, and put $\mathcal{E}=2^{\breve{E}}\setminus\{\emptyset\}$. Consider our updating functions $f_{E}:\mathscr{C}\rightarrow\mathscr{C}$, this time indexed by some $E\in\mathcal{E}$, and let, for all $\mathcal{P}\in\mathscr{C}$, $\overline{f}(\mathcal{P})\coloneqq\overline{\operatorname{CH}(\cup_{E\in\mathcal{E}}f_{E}(\mathcal{P}))}$ and $\underline{f}(\mathcal{P})\coloneqq\cap_{E\in\mathcal{E}}f_{E}(\mathcal{P})$. These maps are well defined because $\overline{f}(\mathcal{P})\in\mathscr{C}$ and we assume $\underline{f}(\mathcal{P})\neq\emptyset$ for all $\mathcal{P}\in\mathscr{C}$. Then, we have that $\underline{f}(\mathcal{P}_{0})\subseteq f_{E}(\mathcal{P}_{0})\subseteq\overline{f}(\mathcal{P}_{0})$, for all $E\in\mathcal{E}$ and all $\mathcal{P}_{0}\in\mathscr{C}$.
Now, fix any $n\in\mathbb{N}$, and define $\bigcirc_{i=1}^{n}f_{E_{i}}\coloneqq f_{E_{n}}\circ\cdots\circ f_{E_{1}}$, where $\{E_{i}\}_{i=1}^{n}\subset\mathcal{E}$. The following is proven in Appendix B.6.
Proposition 5 (Controlling the Limiting Behavior of $\bigcirc_{i=1}^{n}f_{E_{i}}$).
Let $\mathcal{X}$ be compact metric, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Suppose that the following two hypotheses hold,
  • $f_{E}$ is order-preserving (monotone), for all $E\in\mathcal{E}$, that is, for all $\mathcal{P},\mathcal{Q}\in\mathscr{C}$ such that $\mathcal{P}\subseteq\mathcal{Q}$, we have that $f_{E}(\mathcal{P})\subseteq f_{E}(\mathcal{Q})$, for all $E\in\mathcal{E}$;
  • The conditions in Theorem 3 hold for both $\underline{f}$ and $\overline{f}$.
Then, for all $\mathcal{P}_{0}\in\mathscr{C}$, if the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ exists, we have that
$\underline{\mathcal{P}}_{\star}\subseteq\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})\subseteq\overline{\mathcal{P}}_{\star},$
where $\underline{\mathcal{P}}_{\star}\coloneqq\lim_{n\rightarrow\infty}\underline{f}^{n}(\mathcal{P}_{0})$ and $\overline{\mathcal{P}}_{\star}\coloneqq\lim_{n\rightarrow\infty}\overline{f}^{n}(\mathcal{P}_{0})$.
Remark 2.
In many of the credal ML models discussed in Section 3.1, each updater $f_{E}$ is monotone and satisfies a common contraction condition in the Hausdorff metric. In such cases, the induced lower and upper envelope maps $\underline{f}$ and $\overline{f}$ inherit the assumptions of Theorem 3, and the structural and numerical diagnostics in Section 3.1 can be applied directly to the family $(f_{E})_{E\in\mathcal{E}}$.
However, Proposition 5 also shows that this is not sufficient, by itself, to control the limiting behavior of arbitrary compositions: For each $\mathcal{P}_{0}\in\mathscr{C}$, the existence of the Hausdorff limit $\lim_{n\to\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ must still be checked on a case-by-case basis. In other words, we need additional mathematical structure on how the maps $f_{E}$ update the initial credal set $\mathcal{P}_{0}$; Their long-run behavior, when composed, cannot be determined at the level of generality of Proposition 5. Two simple sufficient scenarios are: (a) all $f_{E_{i}}$ are strict Hausdorff contractions with a common Lipschitz constant $L<1$, or (b) the sets $f_{E_{n}}\circ\cdots\circ f_{E_{1}}(\mathcal{P}_{0})$ form a nested chain whose diameters shrink to 0. We prove that they are indeed sufficient in Appendix B.7.
Notice also that, alas, CBDL does not satisfy the order-preserving condition, as (generalized) Bayes’ updating is well known to dilate (Seidenfeld and Wasserman,, 1993). Some mechanisms that satisfy the order-preserving condition can be found in Caprio and Seidenfeld, (2023).
4. Illustration with Finitely Generated Credal Sets
Figure 2. Illustration of FGCSs reaching their fixed points. Starting with a credal set generated by finitely many extreme distributions, we repeatedly apply the updating map $f$. With each iteration, the distance between the parameters of the distributions located at the extreme corners of the set decreases. This contraction effect continues until the successive distances vanish, when the credal set collapses to its fixed point.
We provide a simple illustration of our main result (Theorem 1) using the following set-up. We assume that the data generating process is $X_{1}\dots,X_{n}\sim\operatorname{N}(\theta_{\star},1)$, $\theta_{\star}=0$.
We generate our prior finitely generated credal set $\mathcal{P}_{0}$ by setting $P_{j}(\theta)=\operatorname{N}(\theta\mid\mu_{j},\tau^{2}_{j})$, $j\in\{1,2,3\}$, and $\mathcal{P}_{0}=\text{CH}(\{P_{j}\}_{j=1}^{3})$. Following the notation from Section 3.1.1, we can define the updating procedure of this credal set by $f$, where $f(\mathcal{P})\coloneqq\text{CH}(\varphi(\operatorname{ex}(\mathcal{P}))$, and $\varphi(\operatorname{ex}(\mathcal{P}))$ denotes the elementwise Bayes update with the given evidence $X_{1},\dots,X_{n}$ to each prior $P_{1},P_{2},P_{3}$. As a Gaussian prior on the mean is conjugate to a Gaussian likelihood with known variance, we have a closed form for the posterior finitely generated credal set (Hoff,, 2009), given by $P_{j}(\theta\mid X_{1},\dots X_{n})=\operatorname{N}(\theta\mid\tilde{\mu}_{j},\tilde{\tau}_{j}^{2})$, $j\in\{1,2,3\}$, where $\tilde{\tau}_{j}^{2}=\left(\frac{1}{\tau_{j}^{2}}+n\right)^{-1}$, $\tilde{\mu}_{j}=\tilde{\tau_{j}}^{2}\left(\frac{\mu_{j}}{\tau_{j}^{2}}+\sum_{i=1}^{n}X_{i}\right)$, and $\sum_{i}X_{i}$ is a sufficient statistic for $\theta$.
For the illustration, we set $\mu_{1},\mu_{2},\mu_{3}=(2,3,-2)$ and $\tau_{1}^{2},\tau_{2}^{2},\tau_{3}^{2}=(1,2,2)$. To establish that the credal set $\mathcal{P}$ reaches a fixed point under repeated application of the updating map $f$, it is sufficient to examine only the extreme distributions. Equivalently, instead of tracking full distributions, we only need to verify whether the parameters of the Gaussians $(\mu_{i},\tau_{i}^{2})$ converge to fixed values. If each sequence of parameters stabilizes, then the whole credal set stabilizes.
Figure 2 illustrates the result of repeatedly applying $f$ to the prior finitely generated credal set $\mathcal{P}$. We pack the $t^{\text{th}}$ round of parameters $(\theta_{j,t},\tau^{2}_{j,t})$ together and denote it as $\xi_{t}^{(j)}$. We plot the averaged (over the three extreme points in our FGCSs) consecutive parameter differences $\frac{1}{3}\sum_{j=1}^{3}\|\xi^{(j)}_{t}-\xi^{(j)}_{t-1}\|_{2}^{2}$ on the y-axis and the number of rounds in the x-axis. As illustrated, the difference quickly converges to 0, meaning the credal sets reach their fixed point. We provide the full simulation detail in Appendix C.
5. Discussion
In this work, we made a first stride towards the study of fixed point results for credal sets. In particular, we derived credal sets versions of Kakutani’s, Edelstein’s, and Boyd-Wong’s fixed point theorems. We also discussed how the practitioner can check that the conditions in our results hold, and we studied the fixed point properties of Credal Bayesian Deep Learning (CBDL), a credal set generalization of BDL.
Our findings open up many interesting questions, which will be the topic of future research. First, although intuitive, it remains to be explicitly shown that (Bayesian) posterior consistency / concentration can be framed as special cases of our results. Correctly specifying the likelihood, and then applying Bayes’ rule iteratively, will result in reaching a fixed point being the Dirac distribution at the right parameter. Second, we plan to derive an axiomatization of the desirable properties of an updating mechanism, a goal in line with the recent literature in imprecise opinion pooling (Elkin and Pettigrew,, 2025). Third, notice that, although our examples were based on CBDL, we do not confine our attention to the Bayesian paradigm alone: our results are useful to any credal set learning paradigm. Consider the case where we start with a set of predictors (initializing parameters randomly or having different hyperparameters, as it is done e.g. in deep ensemble) and then, after collecting evidence, we either update the models separately, or “trim” them; Only a few are kept, e.g. those which clear a certain threshold for being the “correct ones”. In the future, we plan to study the structure needed to place on these kinds of threshold-based update rules in addition to the standard learning procedure, to have fixed points, and to be able to reach them. Finally, we plan to inspect how our theorems can be put to use to study the existence of fixed point theorems for lower probabilities in the context of IIPM (Chau et al., 2025a, ), a generalization that take also into account epistemic uncertainty, of the well-studied field of Integral Probability Metrics.
Appendix A Mathematical Background
In this section, we present the preliminary notions that are needed to fully appreciate the results in our paper.
A.1. Compact Metric Space
A metric space $(\mathcal{X},d)$ is compact if and only if every sequence in $\mathcal{X}$ has a convergent subsequence (with respect to $d$) whose limit lies in $\mathcal{X}$ (sequential compactness). Equivalently, $(\mathcal{X},d)$ is compact if and only if it is complete (every Cauchy sequence converges in $\mathcal{X}$) and totally bounded. In typical Machine Learning classification problems, this condition is automatically satisfied when the label space is finite and endowed with the discrete metric. In regression problems, one can enforce compactness by restricting the output space to a compact subset of $\mathbb{R}^{k}$ and using the Euclidean metric, or more generally by replacing the output space with a compactification equipped with a compatible metric.
A.2. Weak Topology
1. General definition of the weak topology.
Let now $\mathcal{X}$ be a normed (or more generally locally convex) space, and let $\mathcal{X}^{\star}$ denote its continuous dual, i.e. the space of all continuous linear functionals $x^{\star}:\mathcal{X}\to\mathbb{K}$, where $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$. The weak topology on $\mathcal{X}^{\star}$, denoted $\sigma(\mathcal{X}^{\star},\mathcal{X})$, is the coarsest topology on $\mathcal{X}^{\star}$ such that, for every $x\in\mathcal{X}$, the evaluation map
$\text{ev}_{x}:\mathcal{X}^{\star}\to\mathbb{K},\qquad\text{ev}_{x}(x^{\star})\coloneqq x^{\star}(x)$
is continuous.
Concretely, a net $(x^{\star}_{\alpha})$ in $\mathcal{X}^{\star}$ converges to $x^{\star}\in\mathcal{X}^{\star}$ in the weak topology if and only if
$x^{\star}_{\alpha}(x)\longrightarrow x^{\star}(x)\qquad\text{for every }x\in\mathcal{X}.$
A typical basic neighborhood of a point $x_{0}^{\star}\in\mathcal{X}^{\star}$ in the weak topology is of the form
$\Bigl\{x^{\star}\in\mathcal{X}^{\star}:\lvert x^{\star}(x_{i})-x_{0}^{\star}(x_{i})\rvert<\varepsilon,i=1,\dots,n\Bigr\},$
for some finite set $\{x_{1},\dots,x_{n}\}\subset X$ and $\varepsilon>0$.
2. Weak topology on a space of probability measures.
Let $\mathcal{Z}$ be a topological space (often assumed to be metric or Polish in probability theory). Denote by $C_{b}(\mathcal{Z})$ the Banach space of bounded, continuous, real-valued functions on $\mathcal{Z}$, equipped with the supremum norm, and by $\Delta_{\mathcal{Z}}$ the set of all Borel probability measures on $\mathcal{Z}$.
Each $\mu\in\Delta_{\mathcal{Z}}$ defines a continuous linear functional on $C_{b}(\mathcal{Z})$ via
$L_{\mu}:C_{b}(\mathcal{Z})\to\mathbb{R},\qquad L_{\mu}(f):=\int_{\mathcal{Z}}f\text{d}\mu.$
Thus we can view $\Delta_{\mathcal{Z}}$ as a subset of the dual space $C_{b}(\mathcal{Z})^{\star}$ through the embedding
$\Delta_{\mathcal{Z}}\hookrightarrow C_{b}(\mathcal{Z})^{\star},\qquad\mu\mapsto L_{\mu}.$
The weak topology $\sigma(C_{b}(\mathcal{Z})^{\star},C_{b}(\mathcal{Z}))$ on $C_{b}(\mathcal{Z})^{\star}$ induces a subspace topology on $\Delta_{\mathcal{Z}}$. This induced topology is exactly the usual weak topology (or narrow topology) on probability measures. Concretely, for a net $(\mu_{\alpha})$ in $\Delta_{\mathcal{Z}}$ and $\mu\in\Delta_{\mathcal{Z}}$,
$\mu_{\alpha}\xrightarrow{\text{weak}^{\star}}\mu\quad\Longleftrightarrow\quad\int_{\mathcal{Z}}f\text{d}\mu_{\alpha}\longrightarrow\int_{\mathcal{Z}}f\text{d}\mu\quad\text{for all }f\in C_{b}(\mathcal{Z}).$
In probability notation, one often writes $\mu_{\alpha}\Rightarrow\mu$ and says that $\mu_{\alpha}$ converges weakly to $\mu$.
A.3. Dominating Measure
Let $(X,\mathcal{F})$ be a measurable space, and call $\mathcal{P}$ a collection of countably additive probability measures on $(X,\mathcal{F})$. Then, a $\sigma$-finite (i.e. countably additive) measure $\mu$ on $(X,\mathcal{F})$ dominates $\mathcal{P}$ if and only if, for all $P\in\mathcal{P}$, $P\ll\mu$. In other words, whenever $\mu(A)=0$, for some $A\in\mathcal{F}$, then $P(A)=0$, for all $P\in\mathcal{P}$.
Whenever a $\sigma$-finite measure $\mu$ dominates $\mathcal{P}$, the Radon-Nikodým theorem guarantees that each $P\in\mathcal{P}$ admits a (measurable) density $p$ with respect to $\mu$, that is, a function $p:X\to[0,\infty)$ such that
$P(A)=\int_{A}p(x)\mu(\mathrm{d}x),\qquad\text{for all }A\in\mathcal{F}.$
Equivalently, we write $p=\frac{\mathrm{d}P}{\mathrm{d}\mu}$ and refer to $p$ as the Radon-Nikodým derivative of $P$ with respect to $\mu$. Thus, a dominating measure $\mu$ allows us to represent all $P\in\mathcal{P}$ by their densities relative to the common reference measure $\mu$.
A.4. Likelihood Kernel
To increase clarity, in this section we keep the notation of Example 3.1.1. There, the function $\ell_{k}$ is a likelihood kernel from the parameter space $(\Theta,\mathcal{T})$ to the evidence space $(\mathcal{Y},\mathcal{F}_{\mathcal{Y}})$. More precisely,
$\ell_{k}:\Theta\times\mathcal{F}_{\mathcal{Y}}\to[0,\infty),\qquad(\theta,E)\mapsto\ell_{k}(E\mid\theta),$
is called a likelihood kernel if the following conditions hold:
  1. (1)
    For each fixed $\theta\in\Theta$, the map
    $E\mapsto\ell_{k}(E\mid\theta)$
    is a (typically probability) measure on $(\mathcal{Y},\mathcal{F}_{\mathcal{Y}})$. In particular, for each $\theta$, $\ell_{k}(\cdot\mid\theta)$ describes the distribution of the evidence, given parameter $\theta$.
  2. (2)
    For each fixed $E\in\mathcal{F}_{\mathcal{Y}}$, the map
    $\theta\mapsto\ell_{k}(E\mid\theta)$
    is $\mathcal{T}$-measurable on $(\Theta,\mathcal{T})$.
Thus, $\ell_{k}$ is a Markov (stochastic) kernel encoding the likelihood: For each $\theta$, $\ell_{k}(\cdot\mid\theta)$ is the conditional law of the evidence, and quantities such as
$L_{k}(E)=\int_{\Theta}\ell_{k}(E\mid\theta)p^{\mathrm{ex}}(\theta)\mu(\mathrm{d}\theta)$
represent the prior predictive (or marginal likelihood) of the event $E$ under the prior $P^{\mathrm{ex}}$.
A.5. Classical Fixed Point Theorems
In this section, we present three classical fixed point theorems that we generalize to the credal set case in the main body of our work.
The first is the general version of Kakutani’s fixed point theorem (Kakutani,, 1941), as stated in Aliprantis and Border, (2006, Corollary 17.55), which is related to Theorem 1.
Theorem 6 (Kakutani-Fan-Glicksberg’s Fixed Point Theorem).
Let $K$ be a nonempty compact convex subset of a locally convex Hausdorff space, and let the set-valued function $f:K\rightarrow K$ have closed graph and nonempty convex values. Then, the set of fixed points of $f$ is compact and nonempty.
The second is Edelstein’s fixed point theorem, as stated in Aliprantis and Border, (2006, Theorem 3.52), which is generalized by Corollary 2.1.
Theorem 7 (Edelstein’s Fixed Point Theorem).
If a function $f:X\rightarrow X$ on a compact metric space $(X,d)$ satisfies $d(f(x),f(y))<d(x,y)$, for all $x\neq y$, then $f$ has a unique fixed point.
Finally, we have Boyd-Wong’s fixed point theorem (Boyd and Wong,, 1969, Theorem 1), which is generalized by Theorem 3.
Theorem 8 (Boyd-Wong’s Fixed Point Theorem).
Let $(X,\rho)$ be a complete metric space. We denote the range of $\rho$ by $R$ and the closure of $R$ by $\overline{R}$, so
$R=\{\rho(x,y):x,y\in X\}.$
Suppose that $T\colon X\to X$ satisfies
(4) $\rho(Tx,Ty)\leq\psi(\rho(x,y)),\qquad x,y\in X,$
where $\psi\colon\overline{R}\to[0,\infty)$ is upper semicontinuous from the right on $\overline{R}$ and satisfies $\psi(t)<t$ for all $t\in\overline{R}\setminus\{0\}$. Then $T$ has a unique fixed point $x_{0}$, and $T^{n}x\to x_{0}$, for each $x\in X$.
Remark 3 (Why Boyd-Wong?).
The reader may ask themselves why, in Theorem 3, we chose to prove a generalization of Boyd-Wong’s fixed point theorem; In other words, one might ask whether the convergence of the orbit $(\mathcal{P}_{n})$ could be established under weaker assumptions.
First, let us discuss $\mathcal{X}$ being Polish. Notice that if $\mathcal{X}$ is not separable, then $\Delta_{\mathcal{X}}$ may not be metrizable, which in turn means that we cannot talk about the Hausdorff metric $d_{H}$ on $\mathscr{C}$, let alone about Hausdorff convergence. Also, if $\mathcal{X}$ is not complete, then there might be an orbit $(\mathcal{P}_{n})$ of some $\mathcal{P}_{0}\in\mathscr{C}$ that has an accumulation point, but does not converge to it.
Regarding $f$, strict contractivity near the fixed point is essential for uniqueness and convergence, and without some form of continuity, fixed point existence may fail.
Now, let us focus on the conditions on $\psi$. If we drop upper semicontinuity, we might still get existence and uniqueness of a fixed point, but we may lose convergence of the orbits. This is because upper semicontinuity ensures that repeated applications of $f$ do not “get stuck” in regions where the contractivity vanishes or oscillates.
Although tricky, the only condition that can be relaxed is that $\psi(t)<t$. Indeed, notice that
  • (i)
    If we let $\psi(t)=t$ on a zero-measure set, as long as $\psi(t)<t$ on a dense subset of $\mathbb{R}_{+}$, and $\psi$ is right-upper semicontinuous, we may retain convergence, but convergence can be slower or conditional, and we may lose uniform convergence rates.
  • (ii)
    Without $\psi(t)<t$ (outside of zero-measure sets), $f$ could be the identity or exhibit limit cycles.
That being said, though, we could ask $\psi(t)\leq q(t)\cdot t$, with $q(t)<1$, for all $t>0$. In this case, convergence should still hold, so to give us a credal set version of Matkowski-type and Wardowski-type fixed point theorems. This extension will be explored properly in future research.
Appendix B Proofs and Additional Results
B.1. Proof for Theorem 1
Endow $\Delta_{\mathcal{X}}$ with any metric $\rho$ that metrizes its weak topology (e.g., the Prokhorov metric). Write $d_{H}$ for the Hausdorff metric on $\mathscr{C}$ induced by $\rho$.
Step 0: Hyperspace compactness.
Since $(\Delta_{\mathcal{X}},\rho)$ is compact metric, the hyperspace $\mathcal{K}(\Delta_{\mathcal{X}})$ of nonempty compact subsets is compact under $d_{H}$. If $P_{n}\to P$ in $d_{H}$ and $x_{n},y_{n}\in P_{n}$ with $x_{n}\to x$, $y_{n}\to y$, then $\lambda x_{n}+(1-\lambda)y_{n}\in P_{n}$ and hence $\lambda x+(1-\lambda)y\in P$. Thus, the subfamily $\mathscr{C}$ of convex sets is closed in $\mathcal{K}(\Delta_{\mathcal{X}})$; Therefore $(\mathscr{C},d_{H})$ is compact.
Step 1: Coordinate support-function embedding.
Let $\mathbf{E}\coloneqq C(\mathcal{X})$ and pick a countable dense set $(f_{k})_{k\geq 1}$ in the unit ball of $\mathbf{E}$. For $P\in\mathscr{C}$ define
$h_{P}(f)\coloneqq\sup_{\mu\in P}\mu(f),\qquad f\in\mathbf{E},$
and set
$J(P)\coloneqq\big(h_{P}(f_{k})\big)_{k\geq 1}\in[-1,1]^{\mathbb{N}}.$
Each coordinate $P\mapsto h_{P}(f_{k})$ is continuous for $d_{H}$, hence $J$ is continuous. Moreover $J$ is affine,
$J\big(\lambda P\oplus(1-\lambda)Q\big)=\lambda J(P)+(1-\lambda)J(Q),\quad\lambda\in[0,1].$
Since $h_{P}$ is $1$-Lipschitz on $\mathbf{E}$ and the family $\{\mu\mapsto\mu(f):f\in\mathbf{E}\}$ separates points in $\Delta_{\mathcal{X}}$, the support function $h_{P}$ determines $P$; As $(f_{k})$ is dense, $J$ is injective. Let
$Z\coloneqq\mathbb{R}^{\mathbb{N}}\text{with the product topology},\qquad K\coloneqq J(\mathscr{C})\subset[-1,1]^{\mathbb{N}}\subset Z.$
Then $K$ is compact (by continuity of $J$ and compactness of $\mathscr{C}$) and convex (by affinity of $J$).
Step 2: Conjugate the update map.
Define $\tilde{f}\coloneqq J\circ f\circ J^{-1}:K\to K$. Because $f$ is $d_{H}$-continuous and $J$ is a homeomorphism onto $K$, $\tilde{f}$ is continuous on $K$.
Step 3: Apply Schauder.
Brouwer-Schauder-Tychonoff fixed point theorem (Aliprantis and Border,, 2006, Theorem 17.56) guarantees a point $h_{*}\in K$ with $\tilde{f}(h_{*})=h_{*}$, since $K$ is compact and convex in the locally convex topological (Hausdorff) vector space $Z$.
Step 4: Pull back to $\mathscr{C}$.
Let $P_{*}\coloneqq J^{-1}(h_{*})$. Then $f(P_{*})=J^{-1}\bigl(\tilde{f}(J(P_{*}))\bigr)=P_{*}$, so $P_{*}$ is a fixed point of $f$. Finally, the collection of fixed points of $f$, $\mathrm{Fix}(f)=\{P\in\mathscr{C}:d_{H}(P,f(P))=0\}$, is closed in compact $\mathscr{C}$, hence compact. $\blacksquare$
B.2. Proof for Proposition 2
Pick any $P^{\text{ex}}\in\text{ex}\mathcal{P}$, any $L_{k}\in\text{ex}\mathcal{L}$, and fix the measurable evidence set $E\subset\mathcal{Y}$. We first show that $g_{k}(P^{\text{ex}})(A)$ is well defined, for all measurable $A\subset\Theta$. Since $\Theta$ is compact and $\ell_{k}(E\mid\cdot)$ is continuous and strictly positive, it attains a minimum $m>0$ and maximum $M>0$ on $\Theta$. Thus, $\ell_{k}(E\mid\cdot)\in\mathscr{L}^{\infty}(\Theta,\mu)$, and the product $\ell_{k}(E\mid\cdot)p^{\text{ex}}(\cdot)\in\mathscr{L}^{1}(\Theta,\mu)$. Therefore,
$L_{k}(E)=\int_{\Theta}\ell_{k}(E\mid\theta)p^{\text{ex}}(\theta)\mu(\text{d}\theta)>0,$
so $g_{k}(P^{\text{ex}})(A)$ is well-defined for all measurable sets $A\subset\Theta$.
We now show continuity in $A$. Set $f_{k}(\theta)\coloneqq\ell_{k}(E\mid\theta)p^{\mathrm{e}x}(\theta)\geq 0$ and define the finite measure
$\nu_{k}(B)\coloneqq\int_{B}f_{k}(\theta)\mu(\text{d}\theta),\qquad B\subset\Theta\text{ measurable}.$
Then
$g_{k}(P^{\mathrm{e}x})(A)=\frac{\nu_{k}(A)}{\nu_{k}(\Theta)}\quad\text{with }\nu_{k}(\Theta)=L_{k}(E)>0.$
If $A_{n}\uparrow A$, by the monotone convergence theorem $\nu_{k}(A_{n})\uparrow\nu_{k}(A)$, hence $g_{k}(P^{\mathrm{e}x})(A_{n})\to g_{k}(P^{\mathrm{e}x})(A)$. If $A_{n}\downarrow A$, since $\nu_{k}$ is finite we have continuity from above: $\nu_{k}(A_{n})\downarrow\nu_{k}(A)$, hence again $g_{k}(P^{\mathrm{e}x})(A_{n})\to g_{k}(P^{\mathrm{e}x})(A)$.
Notice now that, at the beginning of the proof, we picked a generic prior $P^{\text{ex}}\in\text{ex}\mathcal{P}$ and a generic likelihood $L_{k}\in\text{ex}\mathcal{L}$. This means that the functions $A\mapsto g_{k}(P^{\text{ex}})(A)$ are continuous, for all $k$ and all $P^{\text{ex}}$. Then, since $\Delta_{\Theta}$ is locally convex (as a result of $\Theta$ being compact metric), we conclude that $f$ is Hausdorff continuous by Aliprantis and Border, (2006, Theorem 17.37). $\blacksquare$
B.3. Proof of Corollary 2.1
From the proof of Theorem 1, we know that, under our assumptions, $(\mathscr{C},d_{H})$ is a Hausdorff compact metric space. The result, then, follows from Edelstein’s fixed point theorem, Theorem 7 in Appendix A.5. $\blacksquare$
B.4. Proof for Theorem 3
Since $\mathcal{X}$ is a Polish space, the space $\Delta_{\mathcal{X}}$ of countably additive probability measures on $\mathcal{X}$, endowed with the weak topology, is itself a Polish space. Let $\mathscr{F}$ denote the set of all nonempty closed subsets of $\Delta_{\mathcal{X}}$, and notice that $\mathscr{C}\subseteq\mathscr{F}$. We claim that $\mathscr{C}$ is a closed subset of $(\mathscr{F},d_{H})$. Let $\{C_{n}\}\subseteq\mathscr{C}$ be a sequence converging in the Hausdorff metric to some set $C\in\mathscr{F}$. Then, since each $C_{n}$ is nonempty and $\mathscr{F}$ is complete, $C$ is nonempty.
Also, each $C_{n}$ is weak-closed, and limits of such sets under Hausdorff convergence are also closed, hence $C$ is weak-closed.
Finally, each $C_{n}$ is convex, and the limit of convex sets under Hausdorff convergence is convex (this follows from the fact that the convex hull operation is continuous under Hausdorff convergence in a convex metric space).
Thus, $C\in\mathscr{C}$, proving that $\mathscr{C}$ is a closed subset of $\mathscr{F}$. Since $\mathscr{F}$ is complete under $d_{H}$, and $\mathscr{C}$ is a closed subset of $\mathscr{F}$, it follows that $\mathscr{C}$ is complete under $d_{H}$.
Since $(\mathscr{C},d_{H})$ is a compact metric space, then it is also complete. Our claim, then, follows from Boyd-Wong’s fixed point theorem, Theorem 8 in Appendix A.5. $\blacksquare$
Remark 4 (Information about Convergence Rate).
Notice that the only information we get from Theorem 3 about the convergence rate of $(\mathcal{P}_{n})$ is that it is sublinear and governed by $\psi$. For example, if $\psi$ satisfies additional properties—e.g. Lipschitzianity or $\psi(t)=qt$, with $q\in(0,1)$—the rate becomes tractable, as the setting becomes similar to Banach’s. It is also worth mentioning that there is an active literature studying the convergence speed for Boyd-Wong-type contractions under extra assumptions on the structure of $\psi$ (Gupta and Mansotra,, 2023; Gautam and Kaur,, 2023; Singh et al.,, 2022; Das et al.,, 2024).
B.5. Proof for Proposition 4
Fix measurable set $E\subset\mathcal{Y}$.
Hilbert projective metric. For $P,Q\in\Delta_{\Theta}^{++}$ with densities $p,q$ (defined $\mu$-a.e.), set
$\alpha(P,Q)=\operatornamewithlimits{essinf}_{\theta}\frac{p(\theta)}{q(\theta)},\quad\beta(P,Q)=\operatornamewithlimits{esssup}_{\theta}\frac{p(\theta)}{q(\theta)},\quad d_{\mathrm{H}}(P,Q)=\log\Big(\frac{\beta(P,Q)}{\alpha(P,Q)}\Big).$
Let $d_{H}^{\mathrm{H}}$ be the induced Hausdorff metric on nonempty compact subsets of $\Delta_{\Theta}^{++}$.
Lemma 9 (Bayesian tilt is a Hilbert contraction).
Under (III’), for each $k$ the map
$T_{k}(P)=\frac{\ell_{k}(E\mid\cdot)P}{\int_{\Theta}\ell_{k}(E\mid\theta)P(\text{d}\theta)}.$
is a strict contraction in $d_{\mathrm{H}}$ with coefficient
$\tau_{k}=\tanh\Big(\tfrac{1}{4}\log\tfrac{\beta_{k}}{\alpha_{k}}\Big)\in(0,1),$
i.e. $d_{\mathrm{H}}(T_{k}(P_{1}),T_{k}(P_{2}))\leq\tau_{k}d_{\mathrm{H}}(P_{1},P_{2})$.
If $P\in\Delta_{\Theta}^{++}$ has density $p$ with $m\leq p\leq M$, then the density of $T_{k}(P)$ is
$p^{\prime}(\theta)=\frac{\ell_{k}(E\mid\theta)p(\theta)}{\int_{\Theta}\ell_{k}(E\mid\theta)p(\theta)\mu(\text{d}\theta)}.$
Since $\alpha_{k}\leq\ell_{k}(E\mid\theta)\leq\beta_{k}$ and $\int_{\Theta}p\text{d}\mu=1$, we have
$\alpha_{k}\leq\int_{\Theta}\ell_{k}(E\mid\theta)p(\theta)\text{d}\mu\leq\beta_{k}.$
Thus, for $p^{\prime}(\theta)=\dfrac{\ell_{k}(E\mid\theta)p(\theta)}{\int_{\Theta}\ell_{k}(E\mid\theta)p(\theta)\text{d}\mu}$,
$\frac{\alpha_{k}m}{\beta_{k}}\leq p^{\prime}(\theta)\leq\frac{\beta_{k}M}{\alpha_{k}}\quad\mu\text{-a.e. }\theta,$
and therefore $T_{k}(P)\in\Delta_{\Theta}^{++}$.
Proof of Lemma 9.
Multiplication by $\ell_{k}(E\mid\cdot)$ defines a positive linear map $M_{k}:p\mapsto\ell_{k}(E\mid\cdot)p$ on the positive cone of $L^{1}(\mu)$; the normalization $p\mapsto p/\int p\text{d}\mu$ is projective (leaves $d_{\mathrm{H}}$ unchanged). Since $\alpha_{k}\leq\ell_{k}\leq\beta_{k}$ a.e., the projective diameter of $M_{k}$ equals $\mathscr{D}(M_{k})=\log(\beta_{k}/\alpha_{k})$. By the Birkhoff-Bushell theorem,
$d_{\mathrm{H}}(T_{k}(P_{1}),T_{k}(P_{2}))\leq\tanh\Big(\frac{\mathscr{D}(M_{k})}{4}\Big)d_{\mathrm{H}}(P_{1},P_{2})=\tanh\Big(\tfrac{1}{4}\log\tfrac{\beta_{k}}{\alpha_{k}}\Big)d_{\mathrm{H}}(P_{1},P_{2}).$
Proposition 10 (Set-level contraction and $\psi$).
Let $\Phi(A)=\bigcup_{k=1}^{K}T_{k}[A]$ on $\mathcal{K}(\Delta_{\Theta}^{++})$ (nonempty compact subsets of $\Delta_{\Theta}^{++}$), and put $\tau=\max_{k}\tau_{k}<1$. Then,
$d_{H}^{\mathrm{H}}\big(\Phi(A),\Phi(B)\big)\leq\tau d_{H}^{\mathrm{H}}(A,B)\qquad\forall A,B,$
so $\psi(t)=\tau t$ satisfies (i)-(ii) in Corollary 2.1.
Proof of Proposition 10.
Each $T_{k}$ is $\tau_{k}$-Lipschitz in $d_{\mathrm{H}}$ by Lemma 9; hence $A\mapsto T_{k}[A]$ is $\tau_{k}$-Lipschitz in $d_{H}^{\mathrm{H}}$. For finite unions, $d_{H}^{\mathrm{H}}(\cup_{k}A_{k},\cup_{k}B_{k})\leq\max_{k}d_{H}^{\mathrm{H}}(A_{k},B_{k})$, yielding the bound with $\tau=\max_{k}\tau_{k}$. Then $\psi(t)=\tau t$ is continuous, nondecreasing, $\psi(0)=0$, and $\psi(t)<t$ for all $t>0$. ∎
Proposition 4, then, follows immediately from Lemma 9 and Proposition 10. $\blacksquare$
B.6. Proof for Proposition 5
Pick any $\mathcal{P}_{0}\in\mathscr{C}$, and any $E_{1},E_{2}\in\mathcal{E}$, $E_{1}\neq E_{2}$. As we pointed out before, it is easy to see that
$\underline{f}(\mathcal{P}_{0})\subseteq f_{E_{1}}(\mathcal{P}_{0})\subseteq\overline{f}(\mathcal{P}_{0}).$
Now, call $\mathcal{Q}_{1}\coloneqq\underline{f}(\mathcal{P}_{0})$, $\mathcal{Q}_{2}\coloneqq f_{E_{1}}(\mathcal{P}_{0})$, and $\mathcal{Q}_{3}\coloneqq\overline{f}(\mathcal{P}_{0})$. Of course, $\mathcal{Q}_{1}\subseteq\mathcal{Q}_{2}\subseteq\mathcal{Q}_{3}$.
Now, since we assumed that $f_{E}$ is order-preserving for all $E\in\mathcal{E}$, we have that
(5) $f_{E_{2}}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{2})\subseteq f_{E_{2}}(\mathcal{Q}_{3}).$
In addition, by the definitions of $\underline{f}$ and $\overline{f}$, we have that
(6) $\underline{f}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{1})\quad\text{ and }\quad f_{E_{2}}(\mathcal{Q}_{3})\subseteq\overline{f}(\mathcal{Q}_{3}).$
Putting (5) and (6) together, we obtain
$\underline{f}(\underline{f}(\mathcal{P}_{0}))=\underline{f}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{2})=f_{E_{2}}(f_{E_{1}}(\mathcal{P}_{0}))\subseteq\overline{f}(\mathcal{Q}_{3})=\overline{f}(\overline{f}(\mathcal{P}_{0})),$
or, more compactly, $\underline{f}^{2}(\mathcal{P}_{0})\subseteq f_{E_{2}}(f_{E_{1}}(\mathcal{P}_{0}))\subseteq\overline{f}^{2}(\mathcal{P}_{0})$. Clearly, this holds for all $n\geq 2$, which we can write as
$\underline{f}^{n}(\mathcal{P}_{0})\subseteq\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})\subseteq\overline{f}^{n}(\mathcal{P}_{0}).$
Now, assuming that the conditions of Corollary 3 hold for both $\underline{f}$ and $\overline{f}$, and that the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ exists, we can conclude that
$\underline{\mathcal{P}}_{\star}\eqqcolon\lim_{n\rightarrow\infty}\underline{f}^{n}(\mathcal{P}_{0})\subseteq\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})\subseteq\lim_{n\rightarrow\infty}\overline{f}^{n}(\mathcal{P}_{0})\coloneqq\overline{\mathcal{P}}_{\star},$
thus proving the statement. $\blacksquare$
B.7. Two sufficient conditions for existence of the Hausdorff limit
In this section, we show that the two conditions (a) and (b) in Remark 2 are indeed sufficient for the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ to exist. We prove a general version of the results; it is then enough to put $X=\Delta_{\mathcal{X}}$ and $\mathscr{C}$ as our class of credal sets to retrieve our setting.
Let $(X,d)$ be a compact metric space and let $\mathcal{K}(X)$ denote the set of all nonempty compact subsets of $X$, endowed with the Hausdorff metric $d_{H}$. Let $\mathscr{C}\subseteq\mathcal{K}(X)$ and let $(f_{i})_{i\geq 1}$ be a sequence of maps $f_{i}:\mathscr{C}\to\mathscr{C}$. Fix $\mathcal{P}_{0}\in\mathscr{C}$ and define
$\mathcal{P}_{n}\coloneqq f_{n}\circ f_{n-1}\circ\cdots\circ f_{1}(\mathcal{P}_{0}),\qquad n\geq 1.$
Then the following are sufficient for the existence of the Hausdorff limit $\lim_{n\to\infty}\mathcal{P}_{n}$ in $(\mathcal{K}(X),d_{H})$:
  • (a)
    Uniform Hausdorff contraction. There exists a constant $L\in[0,1)$ such that, for all $i\geq 1$ and all $\mathcal{P},\mathcal{Q}\in\mathscr{C}$,
    $d_{H}\bigl(f_{i}(\mathcal{P}),f_{i}(\mathcal{Q})\bigr)\leq Ld_{H}(\mathcal{P},\mathcal{Q}).$
  • (b)
    Nested chain with vanishing diameters. The sequence $(\mathcal{P}_{n})_{n\geq 1}$ is nested and decreasing, i.e. $\mathcal{P}_{n+1}\subseteq\mathcal{P}_{n}$ for all $n\geq 1$, and its diameters satisfy
    $\operatorname{diam}(\mathcal{P}_{n})\coloneqq\sup\{d(x,y):x,y\in\mathcal{P}_{n}\}\longrightarrow 0\quad\text{as }n\to\infty.$
In case (a), the sequence $(\mathcal{P}_{n})$ is Cauchy in $(\mathcal{K}(X),d_{H})$ and hence convergent. In case (b), the sequence $(\mathcal{P}_{n})$ converges in $d_{H}$ to a singleton $\{x_{\star}\}\in\mathcal{K}(X)$.
Proof.
It is well known that if $(X,d)$ is compact, then $(\mathcal{K}(X),d_{H})$ is a compact metric space (hence complete). We use this fact in both parts.
Proof of (a). Assume that there exists $L\in[0,1)$ such that
$d_{H}\bigl(f_{i}(\mathcal{P}),f_{i}(\mathcal{Q})\bigr)\leq Ld_{H}(\mathcal{P},\mathcal{Q})\quad\text{for all }i\geq 1,\mathcal{P},\mathcal{Q}\in\mathscr{C}.$
Define $\mathcal{P}_{n}$ as in the statement and let
$\delta_{n}\coloneqq d_{H}(\mathcal{P}_{n},\mathcal{P}_{n-1}),\qquad n\geq 1.$
Then, for $n\geq 1$,
$\delta_{n+1}=d_{H}\bigl(\mathcal{P}_{n+1},\mathcal{P}_{n}\bigr)=d_{H}\bigl(f_{n+1}(\mathcal{P}_{n}),f_{n+1}(\mathcal{P}_{n-1})\bigr)\leq Ld_{H}(\mathcal{P}_{n},\mathcal{P}_{n-1})=L\,\delta_{n}.$
By induction this yields
$\delta_{n+1}\leq L^{n}\delta_{1}=L^{n}d_{H}(\mathcal{P}_{1},\mathcal{P}_{0}),\qquad n\geq 0.$
Let $m>n\geq 0$. Using the triangle inequality and the above bound,
$d_{H}(\mathcal{P}_{m},\mathcal{P}_{n})\leq\sum_{j=n}^{m-1}d_{H}(\mathcal{P}_{j+1},\mathcal{P}_{j})=\sum_{j=n}^{m-1}\delta_{j+1}\leq\sum_{j=n}^{m-1}L^{j}\delta_{1}=L^{n}\frac{1-L^{m-n}}{1-L}\delta_{1}\leq\frac{L^{n}}{1-L}\delta_{1}.$
Since $L\in[0,1)$, the right-hand side tends to 0 as $n\to\infty$, uniformly in $m>n$. Thus $(\mathcal{P}_{n})$ is a Cauchy sequence in the complete metric space $(\mathcal{K}(X),d_{H})$, and hence there exists $\mathcal{P}_{\star}\in\mathcal{K}(X)$ such that $d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})\to 0$ as $n\to\infty$.
Proof of (b). Assume now that $(\mathcal{P}_{n})$ is nested and decreasing (i.e. $\mathcal{P}_{n+1}\subseteq\mathcal{P}_{n}$ for all $n$) and that $\operatorname{diam}(\mathcal{P}_{n})\to 0$ as $n\to\infty$.
First, we show that $\bigcap_{n=1}^{\infty}\mathcal{P}_{n}$ is a singleton. Pick any sequence $(x_{n})_{n\geq 1}$ such that $x_{n}\in\mathcal{P}_{n}$ for each $n$. For $m>n$ we have $x_{n},x_{m}\in\mathcal{P}_{n}$, because $\mathcal{P}_{m}\subseteq\mathcal{P}_{n}$. Hence
$d(x_{n},x_{m})\leq\operatorname{diam}(\mathcal{P}_{n}).$
As $n\to\infty$ we have $\operatorname{diam}(\mathcal{P}_{n})\to 0$, so $(x_{n})$ is a Cauchy sequence in the complete space $X$, and thus converges to some $x_{\star}\in X$.
Next, for each fixed $n$, the tail $\{x_{m}:m\geq n\}$ is contained in $\mathcal{P}_{n}$, which is closed. Since $x_{m}\to x_{\star}$, it follows that $x_{\star}\in\mathcal{P}_{n}$ for every $n$, so that $x_{\star}\in\bigcap_{n=1}^{\infty}\mathcal{P}_{n}$ and the intersection is nonempty.
To see uniqueness, let $y_{\star}\in\bigcap_{n=1}^{\infty}\mathcal{P}_{n}$ be arbitrary. Then $x_{\star},y_{\star}\in\mathcal{P}_{n}$ for every $n$, hence
$d(x_{\star},y_{\star})\leq\operatorname{diam}(\mathcal{P}_{n})\quad\text{for all }n.$
Letting $n\to\infty$ we obtain $d(x_{\star},y_{\star})=0$, so $x_{\star}=y_{\star}$. Thus $\bigcap_{n=1}^{\infty}\mathcal{P}_{n}=\{x_{\star}\}$ is a singleton.
Finally, we show that $\mathcal{P}_{n}\to\{x_{\star}\}$ in the Hausdorff metric. For each $n$,
$d_{H}(\mathcal{P}_{n},\{x_{\star}\})=\max\left\{\sup_{x\in\mathcal{P}_{n}}d(x,x_{\star}),\sup_{y\in\{x_{\star}\}}\inf_{x\in\mathcal{P}_{n}}d(y,x)\right\}.$
Since $x_{\star}\in\mathcal{P}_{n}$ for all $n$, the second term is zero, and we have
$d_{H}(\mathcal{P}_{n},\{x_{\star}\})=\sup_{x\in\mathcal{P}_{n}}d(x,x_{\star})\leq\operatorname{diam}(\mathcal{P}_{n}).$
As $\operatorname{diam}(\mathcal{P}_{n})\to 0$, it follows that $d_{H}(\mathcal{P}_{n},\{x_{\star}\})\to 0$ as $n\to\infty$. Hence $\mathcal{P}_{n}$ converges in the Hausdorff metric to the singleton $\{x_{\star}\}$. ∎
Appendix C Simulation Details
To reproduce Figure 2, one could run the following script:
Listing 1: Iterative Bayesian Updating Simulation
1# Initialisation
2N = 50
3NUM_ROUNDS = 150
4
5# Data Generating Process
6TRUE_SIGMA, TRUE_MEAN = 2.0, 0.0
7theta = np.random.normal(TRUE_MEAN, TRUE_SIGMA)
8samples = np.random.normal(theta, scale=1.0, size=(N))
9
10# Set up prior parameter sets
11PRIOR_MEANS = [2.0, 3.0, -2.0]
12PRIOR_SIGMAS = [1.0, 2.0, 2.0]
13
14posterior_means_set, posterior_SIGMAs_set = [], []
15for i in range(len(PRIOR_MEANS)):
16 PRIOR_SIGMA = PRIOR_SIGMAS[i]
17 PRIOR_MEAN = PRIOR_MEANS[i]
18
19 posterior_means, posterior_SIGMAs = [], []
20 for t in range(NUM_ROUNDS):
21 posterior_sigma = 1/(1/PRIOR_SIGMA + N/TRUE_SIGMA)
22 posterior_mean = posterior_sigma * (
23 PRIOR_MEAN/PRIOR_SIGMA + N * samples.mean()/TRUE_SIGMA
24 )
25
26 PRIOR_MEAN = posterior_mean
27 PRIOR_SIGMA = posterior_sigma
28
29 posterior_means.append(posterior_mean)
30 posterior_SIGMAs.append(posterior_sigma)
31
32 posterior_means_set.append(np.array(posterior_means))
33 posterior_SIGMAs_set.append(np.array(posterior_SIGMAs))
34
35consecutive_diff_in_mean = 0
36consecutive_diff_in_SIGMA = 0
37for j in range(3):
38 consecutive_diff_in_mean += np.abs(
39 posterior_means_set[j][:-1] - posterior_means_set[0][1:]
40 )
41 consecutive_diff_in_SIGMA += np.abs(
42 posterior_SIGMAs_set[j][:-1] - posterior_SIGMAs_set[0][1:]
43 )
44
45consecutive_diff_in_mean /= 3
46consecutive_diff_in_SIGMA /= 3
47
48plt.figure(figsize=(4.5, 3))
49plt.plot(range(NUM_ROUNDS-1), consecutive_diff_in_SIGMA,
50 label="Consecutive parameter difference", linewidth=3)
51plt.xlabel("Number of Rounds \$t\$")
[…]
53plt.legend()
54plt.show()
Appendix D Additional Result: Finitely Additive Case
The reason why the results in Section 3 may not hold in the finitely additive case is that, if $\mathcal{X}$ is not finite, then the space of finitely additive probabilities may not be metrizable, and so we cannot talk about Hausdorff distance, which played a crucial role in the proofs in Appendix B. On the other hand, the space of countably additive probability measures (under the weak topology) is always metrizable, provided $\mathcal{X}$ is metrizable. These concerns should not worry the practitioner, though, as classical probabilistic ML works with countably additive probability measures.
In this section, we generalize Corollary 3 to the finitely additive case; as we shall see, we need to pay a price for this extension.
As for Section 3, let $\mathcal{X}$ be a measurable space of interest. Call $\Delta^{\text{fa}}_{\mathcal{X}}$ the space of finitely additive probability measures on $\mathcal{X}$, and let $\mathscr{C}^{\text{fa}}\subset 2^{\Delta^{\text{fa}}_{\mathcal{X}}}$ be the space of nonempty weak-closed and convex subsets of $\Delta^{\text{fa}}_{\mathcal{X}}$. Denote by $\mathcal{P}^{\text{fa}}\in\mathscr{C}^{\text{fa}}$ a generic credal set in this extended environment. We first show the following lemma.
Lemma 11 ($\mathcal{X}$ finite $\iff$ $(\mathscr{C}^{\mathrm{fa}},d_{H})$ complete metric).
Endow $\Delta^{\mathrm{fa}}_{\mathcal{X}}$ with the weak topology inherited from $ba(\mathcal{F}_{\mathcal{X}})\cong B_{b}(\mathcal{X})^{\star}$, where $B_{b}(\mathcal{X})$ denotes the bounded $\mathcal{F}_{\mathcal{X}}$-measurable functions with the sup norm. [^2] Here $ba(\mathcal{F}_{\mathcal{X}})$ is the space of bounded finitely additive signed measures on the $\sigma$-algebra $\mathcal{F}_{\mathcal{X}}$. Every $\nu\in ba(\mathcal{F}_{\mathcal{X}})$ defines a continuous linear functional $f\mapsto\int f\text{d}\nu$ on $B_{b}(\mathcal{X})$, and conversely. If $\mathcal{F}_{\mathcal{X}}=2^{\mathcal{X}}$, then $B_{b}(\mathcal{X})=\ell^{\infty}(\mathcal{X})$ and $ba(\mathcal{F}_{\mathcal{X}})=(\ell^{\infty}(\mathcal{X}))^{\star}$. Let $d_{H}$ denote the Hausdorff distance induced by some metric compatible with this weak topology on $\Delta^{\mathrm{fa}}_{\mathcal{X}}$. Then $\mathcal{X}$ is finite, or equivalently, $\mathcal{F}_{\mathcal{X}}$ has finitely many atoms, if and only if, $\big(\mathscr{C}^{\mathrm{fa}},d_{H}\big)$ is complete.
Proof.
Only if. If $\mathcal{X}$ is finite (hence $B_{b}(\mathcal{X})$ is finite dimensional and separable), then the unit ball of $ba(\mathcal{F}_{\mathcal{X}})=B_{b}(\mathcal{X})^{\star}$ is weak compact and metrizable; in particular $\Delta^{\mathrm{fa}}_{\mathcal{X}}$ is a compact metric space in the weak topology. The hyperspace $\mathcal{K}(\Delta^{\mathrm{fa}}_{\mathcal{X}})$ of nonempty compact subsets, endowed with $d_{H}$, is compact (hence complete). The subfamily $\mathscr{C}^{\mathrm{fa}}$ of nonempty weak-closed convex subsets is closed under Hausdorff limits (convexity is preserved), so $\mathscr{C}^{\mathrm{fa}}$ is complete.
If. If $\big(\mathscr{C}^{\mathrm{fa}},d_{H}\big)$ is a complete metric space for the Hausdorff metric induced by a metric compatible with the weak topology on $\Delta^{\mathrm{fa}}_{\mathcal{X}}$, then the weak topology on $\Delta^{\mathrm{fa}}_{\mathcal{X}}$ is metrizable. Hence the weak topology on the unit ball of $ba(\mathcal{F}_{\mathcal{X}})=B_{b}(\mathcal{X})^{\star}$ is metrizable, which by the Grothendieck-Krein-Šmulian criterion (equivalently: weak metrizability of the dual unit ball) implies that $B_{b}(\mathcal{X})$ is separable. This forces $\mathcal{F}_{\mathcal{X}}$ to be finite (in particular, if $\mathcal{F}_{\mathcal{X}}=2^{\mathcal{X}}$, then $\mathcal{X}$ must be finite). Thus $\mathcal{X}$ is finite. ∎
A direct consequence of Lemma 11 is the following, which is a version of Corollary 3 in the finitely additive case.
Theorem 12 (Convergence of $(\mathcal{P}^{\text{fa}}_{n})$).
Let $\mathcal{X}$ be finite, and endow $\Delta^{\text{fa}}_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}^{\text{fa}}\rightarrow\mathscr{C}^{\text{fa}}$ that represents our updating technique. If $f$ is such that
$d_{H}(f(\mathcal{P}^{\text{fa}}),f(\mathcal{Q}^{\text{fa}}))\leq\psi(d_{H}(\mathcal{P}^{\text{fa}},\mathcal{Q}^{\text{fa}})),$
for all $\mathcal{P}^{\text{fa}},\mathcal{Q}^{\text{fa}}\in\mathscr{C}^{\text{fa}}\text{, }\mathcal{P}^{\text{fa}}\neq\mathcal{Q}^{\text{fa}}$, and $\psi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is
  1. (i)
    upper semicontinuous from the right on $\mathbb{R}_{+}$,
  2. (ii)
    such that $\psi(t)<t$, for all $t>0$,
then $f$ has a unique fixed point $\mathcal{P}^{\text{fa}}_{\star}$, and, for each $\mathcal{P}^{\text{fa}}_{0}\in\mathscr{C}^{\text{fa}}$, the orbit $(\mathcal{P}^{\text{fa}}_{n})$ of $\mathcal{P}^{\text{fa}}_{0}$ under $f$ converges to $\mathcal{P}^{\text{fa}}_{\star}$. That is, $\mathcal{P}^{\text{fa}}_{n}\xrightarrow[n\rightarrow\infty]{d_{H}}\mathcal{P}^{\text{fa}}_{\star}$.
Proof.
By Lemma 11, we know that $\mathcal{X}$ finite implies that $(\mathscr{C}^{\text{fa}},d_{H})$ is a complete metric space. The result, then, follows from Boyd-Wong’s fixed point theorem, Theorem 8 in Appendix A.5. ∎
As we can see, in the finitely additive setting, it is not enough for $\mathcal{X}$ to be Polish, as in Theorem 3. To accommodate for finite continuity, we must require that $\mathcal{X}$ is finite. This means that we move from a topological to a set-theoretic structure requirement for $\mathcal{X}$. The latter condition is of course more restrictive than the ones we had in Section 3.